# Communication Optimization for Distributed Training: Architecture, Advances, and Opportunities

### v1

@@ Abstract
The past few years have witnessed the flourishing of large-scale deep neural network models with ever-growing parameter numbers. Training such large-scale models typically requires massive memory and computing resources that exceed those of a single GPU, necessitating distributed training. As GPU performance has rapidly evolved in recent years, computation time has shrunk, thereby increasing the proportion of communication in the overall training time. Therefore, optimizing communication for distributed training has become an urgent issue. In this article, we briefly introduce the general architecture of distributed deep neural network training and analyze relationships among Parallelization Strategy, Collective Communication Library, and Network from the perspective of communication optimization, which forms a three-layer paradigm. We then review current representative research advances with this three-layer paradigm. We find that layers in the current three-layer paradigm are relatively independent, but there is a rich design space for cross-layer collaborative optimization in distributed training scenarios. Therefore, we further advocate a communication-efficient five-layer paradigm underlining opportunities for collaboration designs and look forward to the perspectives of “Vertical”, “Horizontal”, “Intra-Inter” and “Host-Net” collaboration designs. We hope this article can shed some light on future research on communication optimization for distributed training.
@@
Large-scale deep neural network (DNN) models (“large models” for short) have become ubiquitous, and their capabilities have significantly advanced in recent years. The latest prominent large models such as GPT111https://openai.com/gpt-4, LLaMA222https://llama.meta.com/llama2, and GLM333https://open.bigmodel.cn demonstrated unprecedented performance, enabling significant changes in production and lifestyle.
@@
The rapid development of large models inevitably led to a significant increase in the scale of parameters and training data. Due to the limited memory and computing power of a single GPU (for simplicity, we use GPU to signify other AI hardware such as TPU and NPU), training large models with a single GPU has already been a thing of the past. For instance, even merely considering the computation time, training a GPT-3 model with 175 billion parameters on a single Nvidia V100 GPU would require approximately 288 years [1]. Consequently, training large models with multiple GPUs is natural and inevitable, creating a new demand for large-scale, high-performance GPU clusters.
@@
However, enhancing GPU performance and enlarging cluster size do not necessarily lead to linear performance improvement of distributed deep neural network training (“distributed training” for short) systems, as one might naturally expect. This is because the time overhead of distributed training comes not only from computation but also from communication. When the time spent on computation is reduced, the communication time is exposed more and gradually becomes a bottleneck. In fact, the communication tasks account for up to 60% of a DNN training iteration time in Meta’s production environment [2]. Optimizing communication can significantly reduce the overall training time, so there is an urgent need to enhance communication efficiency for distributed training.
Systematically optimizing distributed training is not a straightforward task. Distributed deep neural network training architecture consists of hardware and software parts. The software part includes parallelization strategy, deep learning library, collective communication library (CCL), and network protocols, while the hardware part comprises CPUs, GPUs, RAM, I/O, and network infrastructure. Attaining efficiency in distributed training systems necessitates effective collaboration among various system components[3, 4].
@@
Many components of distributed training architecture are closely related to communication. For instance, parallelization strategy determines communication demand, CCL generates communication traffic, and network affects the efficiency of performing communication tasks. To sum up, there is a “Parallelization Strategy, CCL, and Network” three-layer paradigm of communication optimization for distributed training.
@@
There are many research advances optimizing communication within the current three-layer paradigm. For example, PTD-P [1] uses a novel interleaved pipeline parallelism optimization scheme to overlap communication and computation as much as possible. TACCL [5] generates communication primitive algorithms tailored to specific training tasks and topology to enhance efficiency. TopoOpt [2] leverages the reconfigurability of optical switches to optimize topologies and parallel strategies collaboratively, providing customized topology for efficient communication.
Different from general high-performance computing (HPC) scenarios, distributed training has its own characteristics. For example, communication traffic of distributed training exhibits periodic repetition, and the traffic pattern of each training iteration is relatively consistent [6]. Moreover, most communication tasks in distributed training are pre-determined, which contrasts significantly with traditional tasks where traffic flows arrive stochastically. Meanwhile, the current three-layer paradigm has its shortcomings. Each layer of the paradigm is relatively independent, making it difficult for them to cooperate for communication optimization. Therefore, we advocate a communication-efficient five-layer paradigm, which is detailed in Section IV, aiming to achieve cross-layer collaboration design with (logical) schedulers as middleware. We prospect optimization opportunities from four perspectives: “Vertical”, “Horizontal”, “Intra-Inter”, and “Host-Net” co-design.
@@
The rest of this article is arranged as follows: Section II presents the general architecture of distributed training and the current three-layer paradigm from the perspective of communication optimization. Section III reviews representative research advances on communication optimization for distributed training. Section IV advocates a communication-efficient five-layer paradigm, looking into the cross-layer collaborative design opportunities from four promising research directions. Finally is the conclusion.
@@ II Architecture
This section first provides an overview of the distributed deep neural network training architecture, followed by an in-depth introduction of three components pertinent to communication optimization: parallelization strategy, collective communication library, and network. Finally, the relationships between these three components and communication optimization are elaborated.
@@ II-A Overview
The general architecture of the distributed deep neural network training system is shown in Fig. 3. Top of the architecture is a large deep neural network model to train. As analyzed in Section I, a deep learning model needs distributed training when it is too large to fit into the memory or too slow to train with a single GPU.
Parallelization strategy is an essential part of distributed training deployment, determining how a model is partitioned for distributed training. Commonly used parallelization strategies are detailed in Section II-B.
Deep learning models are often implemented with the deep learning library (such as TensorFlow444https://www.tensorflow.org and PyTorch555https://pytorch.org) , which generates execution task graphs including computing tasks and corresponding collective communication tasks. Communication compression techniques such as gradient quantization are often more relevant to training methods and are not detailed in this article due to the limited space.
The deep learning library often invokes Collective Communication Library to implement collective communication tasks to transmit activations (in forward propagation) or synchronize gradients (in backward propagation) between different GPUs. Commonly used collective communication primitives implemented in CCL are detailed in Section II-C.
The collective communication primitives generate actual communication data traffic, which is injected into the underlying network. A variety of network protocols and topologies for distributed training are detailed in Section II-D. As the end-host of the underlying network, a computing node is a supercomputer equipped with high-performance GPUs and CPUs, a large RAM, and high-speed I/O. These hardware components also need to collaborate efficiently.
@@ II-B Common Parallelization Strategies
Commonly used parallelization strategies include data parallelism, model parallelism (primarily including pipeline parallelism and tensor parallelism), and emerging Mixture-of-Expert (MoE) parallelism, as illustrated by Fig. 3, where each rectangle refers to a GPU node and each round rectangle refers to (part of) a model.
Data parallelism is one of the most commonly used parallelization strategies whose basic idea is to distribute multiple copies of a model to different GPUs. Each GPU receives a subset of training data (mini-batch) during training, and gradients are averaged to update global parameters at the end of each iteration. Model parallelism, including pipeline and tensor parallelism, means splitting the model onto different GPUs. Pipeline parallelism allocates different layers of the model to different GPUs, so there is mainly point-to-point communication between layers (GPUs). Tensor parallelism splits the same layer of the model onto different GPUs and uses distributed matrix computing techniques for collaboration [7], which is a communication-intensive operation. MoE parallelism involves dividing a portion of a model into multiple expert components. Each expert specializes in a specific task domain and is allocated to a certain GPU.
In practice, the above parallelization strategies are often not used independently, but in a hybrid manner. For example, we can first divide the model by layers and split each layer, then distribute the divided model to multiple sets of GPUs, thereby achieving pipeline-tensor-data three-dimensional hybrid parallelism [1]. MoE parallelism is another example of hybrid parallelism. The idea of data parallelism is implicit in MoE parallelism, and the expert model itself can also be split into multiple GPUs for parallel computing.
@@ II-C Common Collective Communication Primitives
Commonly used collective communication primitives implemented in CCL for distributed training are illustrated in Fig. 3, where each rectangle represents a GPU node and each rounded rectangle represents a data chunk.
Broadcast distribute data from a particular node to all other nodes, which can be used in data parallelism and some model parallelisms [8]. All-Gather is a many-to-many collective communication primitive where data from different nodes are distributed to all nodes. All-to-All transmit data among various nodes, such as data distribution in MoE parallelism [9, 10]. All-Reduce is a sum operation of the corresponding data chunk of each node, as shown in Fig. 3(d), where each grey rounded rectangle represents the aggregated results of the data chunks at the corresponding position of each GPU node. Common All-Reduce scenarios include data parallelism and some model parallelism such as Megatron-lm [7].
@@ II-D Underlying Network
Protocols and topologies are two main factors that affect network performance. Besides general TCP/IP protocol, distributed training often uses RDMA (such as RoCE or Infiniband) for less overhead and higher bandwidth.
The traffic flow within the network is also closely related to the network topology. Common network topologies used for distributed training include Fat-tree and its variants, Torus, as well as Ring and Full-Mesh topologies. These topologies can be combined according to practical needs. For example, the NVLink topology of Nvidia’s DGX-1666https://images.nvidia.com/content/pdf/dgx1-system-architecture-whitepaper1.pdf is a combination of Ring and Full-mesh.
II-E Communication Paradigm in the Architecture
Parallelization strategy, CCL, and network form a three-layer paradigm of communication optimization for distributed training, as shown by the colored rectangles in Fig. 3. All of them directly or indirectly affect the communication performance of the training process. Different parallelization strategies determine the primarily used collective communication primitives in task graphs, which affects the traffic pattern. Various implementations of CCL directly affect the actual network traffic. The same traffic demand often exhibits different performance under different network infrastructures. The topologies designed for distributed training are closely related to the algorithm implementation of collective communication primitives. For example, the execution process of the Ring-based All-Reduce algorithm results in a ring communication mode. That is, each node transmits data to its logical neighbors while all the nodes in the communicator (a communicator is a set of nodes used to implement a collective communication task) form a ring. Nvidia’s DGX-16 and Google’s Torus [4] topology both contain numerous ring structures, which are well-suited to satisfy the communication needs of Ring-based All-Reduce.
However, layers are relatively independent in the current three-layer communication paradigm. Although some methods introduced in Section III involve cross-layer design, collaborative design has not yet become mainstream in general, which limits the end-to-end training performance [6]. Therefore, we advocate a communication-efficient five-layer paradigm with additional (logical) schedulers (dashed boxes in Fig. 3) over the current architecture. The inter-layer schedulers imply collaborative design, which is elaborated in Section IV.
@@ III Overview of Recent Advances
Communication optimization in distributed training encompasses numerous aspects. We delve into three of the most pertinent elements: parallelization strategy, CCL, and network, which make up the current three-layer communication paradigm. We use exposed communication to denote the communication time that cannot overlap with computation time. In this section, our focus is confined to the contributions of a specific study pertaining to communication optimization, that is, how to reduce exposed communication in distributed training. Table I provides an overview of the research advances with some primary concerns.
Layer |
Main focus |
How to reduce exposed communication |
Cross-layer co-design |
Time for decision |
Implementation |
Method |
Beneficial effect |
Megatron-lm [7] |
Model Parallelism |
Efficient intra-layer model parallel strategy |
An efficient tensor partitioning strategy |
Reducing the amount of traffic by removing a synchronization point |
Offline |
Insert a few communication operations in naive pytorch |
AlpaComm [8] |
Addressing “cross-mesh resharding” problem in model parallelism |
Broadcast based resharding method and eager 1F1B pipeline schedule |
Reducing inter-node traffic and overlap communication with computation |
With CCL |
On/Offline |
900 lines of C++ and 2.5K lines of python |
PTD-P [1] |
Hybrid Parallelism |
Combination of pipeline, tensor and data parallelisms |
A interleaved pipelining schedule |
Overlapping communication and computation in pipeline parallelism |
Offline |
An extension to Megatron-LM codebase |
Lina [9] |
MoE Parallelism |
All-to-All bottleneck in MoE parallelism |
Priority scheduling All-to-All with All-Reduce micro-ops |
Overlapping All-Reduce communication with computation |
With CCL |
Offline |
Based on DeepSpeed MoE with 7500 lines of code |
Janus [10] |
A data-centric paradigm for MoE parallelism |
Move experts to reduce All-to-All traffic when benificial and a provident prefetch strategy |
Reducing All-to-All traffic and overlap communication with computation |
On/Offline |
A plugin of 4.3K lines of code in PyTorch |
NCCL |
Collective Communi- cation Library (CCL) |
Multi-GPU and multi-node communication primitives optimized for NVIDIA hardware |
Select communication primitive algorithms such as DBTree and Ring for All-Reduce, etc. |
Mainly accelerating All-Reduce communication |
For particular hardware |
On/Offline |
Integrated in Deep learning frameworks such as PyTorch and TensorFlow. |
Blink [11] |
Synthesize collective communication algorithms for underlying topology |
Packing spanning trees with MWU and ILP algorithms |
Overlap flows within primitives to promote link utilization |
For topology |
Online |
Seamlessly plugged into pytorch and tensorflow |
SCCL [12] |
Synthesize collective communication algorithms for underlying topology |
Synthesizing pareto-optimal algorithms with SMT encoding |
Overlap flows within primitives to promote link utilization |
For topology |
Offline |
Evaluated on different 8-GPU systems such as Nvidia DXG-1 and Gigabyte Z52 |
TACCL [5] |
Synthesize collective communication algorithms for topology and training task |
Routing (MILP), ordering (Greedy algorithm) and scheduling (MILP) with communication sketches |
Overlap flows within primitives to promote link utilization |
For topology |
Offline |
Extension of NCCL with TACCL runtime as interpreter |
SYNDI-CATE [13] |
Co-optimize collective communication algorithms with runtime schedule |
MCMC Search with termination |
Overlap primitives to promote link utilization |
With execution plan |
On/Offline |
UCC library interface (pytorch integrated) |
TPUv4 [4] |
Network |
Google’s third supercomputer specific for ML |
Architectural |
Up to 4096 chips connected by OCSs |
TopoOpt [2] |
Co-optimizing computation, communication and network topology |
A group theory-inspired algorithm called TotientPerms |
The best network topology and routing plan, together with a parallelization strategy |
With parallelization strategy |
Offline |
A 12-node prototype with FlexFlow’s training engine and modification to NCCL |

### v2

@@ Abstract
The past few years have witnessed the flourishing of large-scale deep neural network models with ever-growing parameter numbers. Training such large-scale models typically requires massive memory and computing resources, necessitating distributed training. As GPU performance has rapidly evolved in recent years, computation time has shrunk, making communication a larger portion of the overall training time. Consequently, optimizing communication for distributed training has become crucial. In this article, we briefly introduce the general architecture of distributed deep neural network training and analyze relationships among Parallelization Strategy, Collective Communication Library, and Network from the perspective of communication optimization, which forms a three-layer paradigm. We then review current representative research advances within this three-layer paradigm. We find that layers in the current three-layer paradigm are relatively independent and there is a rich design space for cross-layer collaborative optimization in distributed training scenarios. Therefore, we advocate “Vertical” and “Horizontal” co-designs which extend the three-layer paradigm to a five-layer paradigm. We also advocate “Intra-Inter” and “Host-Net” co-designs to further utilize the potential of heterogeneous resources. We hope this article can shed some light on future research on communication optimization for distributed training.
@@
Large-scale deep neural network (DNN) models (“large models” for short) have become ubiquitous, and their capabilities have significantly advanced in recent years. The latest prominent large models such as GPT111https://openai.com/gpt-4, LLaMA222https://llama.meta.com, and GLM333https://open.bigmodel.cn demonstrated unprecedented performance, enabling significant changes in production and lifestyle.
@@
The rapid development of large models leads to a significant increase in the scale of parameters and training data. Due to the limited memory and computing power of a single GPU, training large models with a single GPU has already been a thing of the past. For instance, training a GPT-3 model with 175 billion parameters on a single Nvidia V100 GPU would require approximately 288 years [1]. Consequently, training large models with multiple GPUs is natural and inevitable. This creates a new demand for large-scale, high-performance GPU clusters to accelerate the model training process.
@@
However, simply enhancing GPU performance and enlarging cluster size do not necessarily lead to linear performance improvement of distributed deep neural network training (“distributed training” for short) systems, as one might naturally expect. This is because the time overhead of distributed training comes not only from computation but also from communication. When the time spent on computation is reduced, the communication time is exposed more and gradually becomes a bottleneck. The communication overhead can account for up to 60% of a DNN training iteration time in Meta’s production environment [2]. Therefore, there is an urgent need to enhance communication efficiency for distributed training.
@@
Many components of distributed training systems[3, 4] are closely related to communication. For instance, parallelization strategy determines communication demand, collective communication library (CCL) generates communication traffic, and network affects the efficiency of performing communication tasks. To sum up, there is a Parallelization Strategy, CCL, and Network three-layer paradigm of communication optimization for distributed training.
@@
There are many research advances optimizing communication for distributed training in the current three-layer paradigm. For example, PTD-P [1] uses a novel interleaved pipeline parallelism optimization scheme to overlap communication and computation as much as possible. TACCL [5] generates communication primitive algorithms tailored to specific training tasks and topology to enhance efficiency. TopoOpt [2] leverages the reconfigurability of optical switches to optimize topologies and parallel strategies collaboratively, providing customized topology for efficient communication.
Unlike general high-performance computing scenarios, distributed training has distinct characteristics. For instance, the communication traffic of each iteration is typically fixed [6] and most of the communication tasks are pre-determined, which contrasts with the stochastic traffic in traditional scenarios. However, components in the current communication paradigm often function independently, hindering collaborative optimization for communication. In Section IV, we explore research opportunities from a collaborative design perspective to reduce job completion time (JCT) and enhance the efficiency of distributed training systems.
@@
The rest of this article is arranged as follows: Section II presents the general architecture of distributed training and the current three-layer paradigm from the perspective of communication optimization. Section III reviews representative research advances. Section IV illustrates the necessity of collaborative design through a case study and provides an outlook on research opportunities. Finally is the conclusion.
@@ II Architecture
This section first provides an overview of the distributed training architecture from the perspective of communication, followed by an in-depth introduction of three components pertinent to communication optimization: parallelization strategy, CCL, and network. Finally, we summarize the three components as a three-layer paradigm of communication optimization and analyze their impacts on communication efficiency.
@@ II-A Overview
From the perspective of communication, the general architecture of the distributed training system is shown in Fig. 1. Top of the architecture is a large deep neural network model that requires distributed training.
Parallelization strategy is an essential part of distributed training deployment which determines how a model is partitioned for distributed training. Commonly used parallelization strategies are detailed in Section II-B.
Deep learning models are often implemented with the deep learning library (such as TensorFlow444https://www.tensorflow.org and PyTorch555https://pytorch.org) , which generates execution task graphs including computing tasks and corresponding collective communication tasks.
The deep learning library often invokes collective communication library to implement collective communication primitives to transmit activations (in forward propagation) or synchronize gradients (in backward propagation) between different GPUs. CCL often gathers several GPU nodes as a communicator and schedules data traffic to finish a collective communication task. Commonly used collective communication primitives are detailed in Section II-C.
The actual communication data traffic generated by CCL is injected into the underlying network. A variety of network protocols and topologies for distributed training are detailed in Section II-D.
@@
In order to reduce communication overhead, various communication compression methods are also commonly used in practice but often sacrifice accuracy or memory. Communication compression methods are often associated with deep learning libraries and are out of the scope of this article due to space constraints.
@@ II-B Common Parallelization Strategies
Fig. 2 shows four commonly used parallelization strategies that include data parallelism, model parallelism (includes pipeline and tensor parallelism), and emerging Mixture-of-Expert (MoE) parallelism.
Data parallelism is one of the most commonly used parallelization strategies that distribute a subset of training data (mini-batch) to multiple copies of a model on different GPUs. Model parallelism, including pipeline and tensor parallelism, means splitting the model onto different GPUs. Pipeline parallelism allocates different layers of the model to different GPUs, so there is mainly point-to-point communication between layers (GPUs). Tensor parallelism splits the same layer of the model onto different GPUs and uses distributed matrix computing techniques for collaboration [7], which is a communication-intensive operation. MoE parallelism involves dividing a portion of a model into multiple expert components. Each expert specializes in a specific task domain and is allocated to a certain GPU.
In practice, the above parallelization strategies are often not used independently, but in a hybrid manner. For example, we can first divide the model by layers and split each layer, then distribute the divided model to multiple sets of GPUs, thereby achieving pipeline-tensor-data three-dimensional hybrid parallelism [1]. MoE parallelism is another example of hybrid parallelism. The idea of data parallelism is implicit in MoE parallelism, and the expert model itself can also be split onto multiple GPUs for parallel computation.
@@ II-C Common Collective Communication Primitives
Commonly used collective communication primitives implemented in CCL are illustrated in Fig. 3.
Broadcast distributes data from a particular node to all other nodes, which can be used in data parallelism and some model parallelism [8]. All-Gather is a many-to-many collective communication primitive where data from different nodes are distributed to all nodes. All-to-All transmits data among various nodes, such as data distribution in MoE parallelism [9, 10]. All-Reduce is a sum operation of the corresponding data chunk of each node, as shown in Fig. 3(d), where each grey rounded rectangle represents the aggregated results of the data chunks at the corresponding position of each GPU node. Common All-Reduce scenarios include data parallelism and some model parallelism such as Megatron-lm [7].
@@ II-D Underlying Network
Protocols and topologies are two main factors that affect network performance. Besides general TCP/IP protocol, distributed training often uses RDMA (Remote Direct Memory Access), such as RoCE or Infiniband, for less overhead and higher bandwidth.
The traffic flow within the network is also closely related to the network topology. Common network topologies used for distributed training include Fat-tree and its variants, Torus, as well as Ring and Full-Mesh topologies. These topologies can be combined according to actual needs. For example, the NVLink topology of Nvidia’s DGX-1666https://images.nvidia.com/content/pdf/dgx1-system-architecture-whitepaper1.pdf is a combination of Ring and Full-mesh.
II-E Three-Layer Paradigm of Communication Optimization
Parallelization strategy, CCL, and network form a three-layer paradigm of communication optimization for distributed training, as shown by the colored rectangles in Fig. 1. All of them directly or indirectly affect the communication performance of the training process. Different parallelization strategies determine the primarily used collective communication primitives in task graphs, which affect the traffic pattern. Various implementations of a communication primitive in CCL directly affect the network traffic. The same traffic demand often exhibits different performance under different network infrastructures. The topologies designed for distributed training are closely related to the algorithm implementation of collective communication primitives. For example, the execution process of the Ring-based All-Reduce algorithm results in a ring communication mode. That is, each node transmits data to its logical neighbors while all the nodes in the communicator (a communicator is a set of nodes used to perform a collective communication task) form a ring. Nvidia’s DGX-16 and Google’s Torus [4] topology both contain numerous ring structures, which are well-suited to satisfy the communication needs of Ring-based All-Reduce.
In Section III, we will introduce the recent advances of communication optimization in each of the three layers. We find that inter-layer collaboration is still limited and is a direction worth exploring. Therefore, we advocate a new five-layer paradigm in Section IV, which includes inter-layer schedulers to achieve more efficient communication.
@@ III Overview of Recent Advances
| Specific works | Layer | Focus | How to reduce exposed communication | Time for decision | Scala- bility | Beneficial effect | Deployability |
| Megatron-lm [7] | Para. | Intra-layer model parallelism | Reduce traffic by removing a synchronization point | Offline | High | 74% linear scaling on 512 GPUs | Integrated in PyTorch |
| Alpa-Comm [8] | “Cross-mesh resharding” problem | Reduce inter-node traffic and overlap with computation | Offline & Online | Mid | 1.1x training speedup on GPT | 900 lines of C++ and 2.5K lines of python | |
| PTD-P [1] | Combination of different parallelism | Overlap with computation by a interleaved pipeline | Offline | High | 502 petaFLOP/s on 3072 GPUs | Integrated in PyTorch | |
| Lina [9] | All-to-All bottleneck in MoE parallelism | Overlap All-Reduce comm. with computation | Offline | Mid | Up to 1.73x training speedup | Integrated in PyTorch | |
| Janus [10] | Data-centric paradigm for MoE parallelism | Reduce All-to-All traffic and overlap with computation | Offline & Online | High | Up to 2.4x training speedup | Integrated in PyTorch | |
| NCCL | CCL | Comm. primitive implementation | Select efficient communication primitive algorithms | Offline & Online | High | Widely used infrastructure | Integrated in PyTorch and TensorFlow |
| Blink [11] | Comm. primitive optimization for underlying topology | Overlap flows by fine-grained scheduling to improve link utilization and shorten primitive completion time | Online | Low | Up to 1.4x training speedup | Integrated in PyTorch and TensorFlow | |
| SCCL [12] | Offline | Low | 1.14x to 2.2x training speedup | Integrated in MSCCL (Microsoft CCL) | |||
| TACCL [5] | Synthesizing comm. primitive with human sketch | Offline | Mid | Up to 2.36x training speedup on BERT | Integrated in PyTorch | ||
| SYNDI-CATE [13] | Co-optimization of runtime schedule and execution | Overlap primitives to promote link utilization | Offline & Online | Mid | 1.21x to 1.74x training speedup | Integrated in PyTorch | |
| TPUv4 [4] | Net. | Supercomputer for machine learning | Efficient topology for distributed machine learning | Offline | High | Nearly 10x faster overall | Up to 4096 chips connected by OCSs |
| TopoOpt [2] | Co-optimization for topology and parallelization strategy | Match topology with parallelization strategy | Offline | High | Up to 3.4x training speedup | With FlexFlow’s training engine and modified NCCL |
@@
Communication optimization for distributed training encompasses numerous aspects. We delve into three of the most pertinent elements: parallelization strategy, CCL, and network. In this section, our focus is confined to the contributions of a specific study pertaining to communication optimization, that is, how to reduce exposed communication (the communication time that cannot overlap with computation time) in distributed training. Table I provides an overview of the research advances with some primary concerns. In the table, we use Para., CCL and Net. to represent Parallelization Strategy, Collective Communication Library and Network respectively. We evaluate the studies from many aspects such as time for decision, scalability, and deployability. Time for decision indicates whether it plays a role during training (online) or before training (offline), which is crucial for practical applications. Scalability refers to the ability to be applied to large scenarios. We also show representative performance metrics under specific experimental settings in beneficial effect.